# Weakly Supervised Co-training with Swapping Assignments for Semantic Segmentation

### v1

@@ Appendix B Further Analysis
Method bkg plane bike bird boat bottle bus car cat chair cow AdvCAM [34] CVPR21 90.0 79.8 34.1 82.6 63.3 70.5 89.4 76.0 87.3 31.4 81.3 MCT [65] CVPR22 91.9 78.3 39.5 89.9 55.9 76.7 81.8 79.0 90.7 32.6 87.1 ToCo [53] CVPR23 91.1 80.6 48.7 68.6 45.4 79.6 87.4 83.3 89.9 35.8 84.7 Xu et al. [66] CVPR23 92.4 84.7 42.2 85.5 64.1 77.4 86.6 82.2 88.7 32.7 83.8 BECO [50] CVPR23 91.1 81.8 33.6 87.0 63.2 76.1 92.3 87.9 90.9 39.0 90.2 CoSA* (Ours) 93.1 85.5 48.5 88.7 70.0 77.6 90.4 86.4 90.3 47.2 88.7 Method table dog horse mbike person plant sheep sofa train tv mIoU AdvCAM [34] CVPR21 33.1 82.5 80.8 74.0 72.9 50.3 82.3 42.2 74.1 52.9 68.1 MCT [65] CVPR22 57.2 87.0 84.6 77.4 79.2 55.1 89.2 47.2 70.4 58.8 71.9 ToCo [53] CVPR23 60.5 83.7 83.7 76.8 83.0 56.6 87.9 43.5 60.5 63.1 71.1 Xu et al. [66] CVPR23 59.0 82.4 80.9 76.1 81.4 48.0 88.2 46.4 70.2 62.5 72.2 BECO [50] CVPR23 41.6 85.9 86.3 81.8 76.7 56.7 89.5 54.7 64.3 60.6 72.9 CoSA* (Ours) 54.1 87.3 87.1 79.6 85.6 53.2 89.9 71.9 65.1 63.4 76.4
@@
Class
MCT[65] (CVPR22)
Xu et al. [66] (CVPR23)
ToCo[53] (CVPR23)
CoSA (Ours)
Class
MCT[65] (CVPR22)
Xu et al. [66] (CVPR23)
ToCo[53] (CVPR23)
CoSA (Ours)
background
82.4
85.3
68.5
84.0
wine glass
27.0
33.8
20.6
42.1
person
62.6
72.9
28.1
70.3
cup
29.0
35.8
26.0
33.1
bicycle
47.4
49.8
39.7
52.4
fork
23.4
20.0
7.6
24.2
car
47.2
43.8
38.9
54.3
knife
12.0
12.6
18.4
32.9
motorcycle
63.7
66.2
55.1
71.9
spoon
6.6
6.7
3.0
9.0
airplane
64.7
69.2
62.1
74.0
bowl
22.4
23.7
19.8
22.8
bus
64.5
69.1
39.0
77.2
banana
63.2
64.4
71.5
69.3
train
64.5
63.7
48.7
60.0
apple
44.4
50.8
55.5
61.3
truck
44.8
43.4
37.3
55.4
sandwich
39.7
47.0
41.2
48.3
boat
42.3
42.3
49.1
52.1
orange
63.0
64.6
70.6
69.2
traffic light
49.9
49.3
47.3
55.1
broccoli
51.2
50.6
56.7
52.8
fire hydrant
73.2
74.9
69.6
78.8
carrot
40.0
38.6
46.4
59.4
stop sign
76.6
77.3
70.1
82.2
hot dog
53.0
54.0
60.1
59.9
parking meter
64.4
67.0
67.9
71.5
pizza
62.2
64.1
54.9
56.5
bench
32.8
34.1
43.9
50.2
donut
55.7
59.7
61.1
71.1
bird
62.6
63.1
58.6
65.4
cake
47.9
50.6
42.5
57.0
cat
78.2
76.2
74.0
79.8
chair
22.8
24.5
24.1
33.8
dog
68.2
70.6
64.0
72.8
couch
35.0
40.0
44.2
58.1
horse
65.8
67.1
66.1
71.4
potted plant
13.5
13.0
27.4
23.5
sheep
70.1
70.8
67.9
74.3
bed
48.6
53.7
54.0
61.5
cow
68.3
71.2
69.0
74.0
dining table
12.9
19.2
25.6
29.2
elephant
81.6
82.2
79.7
81.9
toilet
63.1
66.6
62.0
69.7
bear
80.1
79.6
76.8
85.3
tv
47.9
50.8
49.1
53.2
zebra
83.0
82.8
77.5
76.3
laptop
49.5
55.4
55.7
63.9
giraffe
76.9
76.7
66.1
68.5
mouse
13.4
14.4
8.6
16.4
backpack
14.6
17.5
20.3
28.6
remote
41.9
47.1
56.6
49.1
umbrella
61.7
66.9
70.9
73.4
keyboard
49.8
57.2
41.8
49.6
handbag
4.5
5.8
8.1
11.9
cellphone
54.1
54.9
58.5
66.2
tie
25.2
31.4
33.4
47.7
microwave
38.0
46.1
55.5
53.2
suitcase
46.8
51.4
55.3
63.8
oven
29.9
35.3
36.2
49.2
frisbee
43.8
54.1
39.6
63.1
toaster
0.0
2.0
0.0
0.0
skis
12.8
13.0
4.0
22.5
sink
28.0
36.1
19.0
41.9
snowboard
31.4
30.3
15.5
40.5
refrigerator
40.1
52.7
51.9
62.0
sports ball
9.2
36.1
11.0
33.1
book
32.2
34.8
31.5
37.8
kite
26.3
47.5
40.7
59.9
clock
43.2
51.5
32.9
55.2
baseball bat
0.9
7.0
1.8
3.8
vase
22.6
25.8
33.3
33.8
baseball glove
0.7
10.4
17.6
37.9
scissors
32.9
30.7
49.8
54.7
skateboard
7.8
15.2
13.3
12.5
teddy bear
61.9
61.4
67.5
69.3
surfboard
46.5
51.5
21.5
16.5
hair drier
0.0
1.3
10.0
0.3
tennis racket
1.4
26.4
6.8
7.2
toothbrush
12.2
19.0
29.3
39.3
bottle
31.1
37.1
25.7
35.1
mIoU
42.0
45.9
42.4
51.1
Table 13: Per-class Segmentation on COCO val Split. Comparison of per-class segmentation results on the COCO 2014 val set. CoSA is compared with MCTformer, Xu et al. and ToCo. Best results are in bold.
Appendix C Further Implementation Details

### v2

@@ 0.C.3 Per-class Segmentation Comparisons
Class MCT[65] (CVPR22) Xu et al. [66] (CVPR23) ToCo[53] (CVPR23) CoSA* (Ours) Class MCT[65] (CVPR22) Xu et al. [66] (CVPR23) ToCo[53] (CVPR23) CoSA* (Ours)
background 82.4 85.3 68.5 84.0 wine glass 27.0 33.8 20.6 42.1
person 62.6 72.9 28.1 70.3 cup 29.0 35.8 26.0 33.1
bicycle 47.4 49.8 39.7 52.4 fork 23.4 20.0 7.6 24.2
car 47.2 43.8 38.9 54.3 knife 12.0 12.6 18.4 32.9
motorcycle 63.7 66.2 55.1 71.9 spoon 6.6 6.7 3.0 9.0
airplane 64.7 69.2 62.1 74.0 bowl 22.4 23.7 19.8 22.8
bus 64.5 69.1 39.0 77.2 banana 63.2 64.4 71.5 69.3
train 64.5 63.7 48.7 60.0 apple 44.4 50.8 55.5 61.3
truck 44.8 43.4 37.3 55.4 sandwich 39.7 47.0 41.2 48.3
boat 42.3 42.3 49.1 52.1 orange 63.0 64.6 70.6 69.2
traffic light 49.9 49.3 47.3 55.1 broccoli 51.2 50.6 56.7 52.8
fire hydrant 73.2 74.9 69.6 78.8 carrot 40.0 38.6 46.4 59.4
stop sign 76.6 77.3 70.1 82.2 hot dog 53.0 54.0 60.1 59.9
park meter 64.4 67.0 67.9 71.5 pizza 62.2 64.1 54.9 56.5
bench 32.8 34.1 43.9 50.2 donut 55.7 59.7 61.1 71.1
bird 62.6 63.1 58.6 65.4 cake 47.9 50.6 42.5 57.0
cat 78.2 76.2 74.0 79.8 chair 22.8 24.5 24.1 33.8
dog 68.2 70.6 64.0 72.8 couch 35.0 40.0 44.2 58.1
horse 65.8 67.1 66.1 71.4 plant 13.5 13.0 27.4 23.5
sheep 70.1 70.8 67.9 74.3 bed 48.6 53.7 54.0 61.5
cow 68.3 71.2 69.0 74.0 table 12.9 19.2 25.6 29.2
elephant 81.6 82.2 79.7 81.9 toilet 63.1 66.6 62.0 69.7
bear 80.1 79.6 76.8 85.3 tv 47.9 50.8 49.1 53.2
zebra 83.0 82.8 77.5 76.3 laptop 49.5 55.4 55.7 63.9
giraffe 76.9 76.7 66.1 68.5 mouse 13.4 14.4 8.6 16.4
backpack 14.6 17.5 20.3 28.6 remote 41.9 47.1 56.6 49.1
umbrella 61.7 66.9 70.9 73.4 keyboard 49.8 57.2 41.8 49.6
handbag 4.5 5.8 8.1 11.9 cellphone 54.1 54.9 58.5 66.2
tie 25.2 31.4 33.4 47.7 microwave 38.0 46.1 55.5 53.2
suitcase 46.8 51.4 55.3 63.8 oven 29.9 35.3 36.2 49.2
frisbee 43.8 54.1 39.6 63.1 toaster 0.0 2.0 0.0 0.0
skis 12.8 13.0 4.0 22.5 sink 28.0 36.1 19.0 41.9
snowboard 31.4 30.3 15.5 40.5 refrigerator 40.1 52.7 51.9 62.0
sports ball 9.2 36.1 11.0 33.1 book 32.2 34.8 31.5 37.8
kite 26.3 47.5 40.7 59.9 clock 43.2 51.5 32.9 55.2
baseball bat 0.9 7.0 1.8 3.8 vase 22.6 25.8 33.3 33.8
glove 0.7 10.4 17.6 37.9 scissors 32.9 30.7 49.8 54.7
skateboard 7.8 15.2 13.3 12.5 teddy bear 61.9 61.4 67.5 69.3
surfboard 46.5 51.5 21.5 16.5 hair drier 0.0 1.3 10.0 0.3
racket 1.4 26.4 6.8 7.2 toothbrush 12.2 19.0 29.3 39.3
bottle 31.1 37.1 25.7 35.1 mIoU 42.0 45.9 42.4 51.1
Table 11: Per-class Segmentation Results on COCO. Comparison of per-class segmentation results on COCO val. CoSA is compared with MCT, Xu et al. and ToCo. Best results are in bold.
0.C.4 Further Qualitative Comparisons
@@
Transformation Description Parameter Setting
RandomRescale Rescale the image by $r$ times, $r$ randomly sampled from $r\sim U(r_{min},r_{max})$. $r_{min}=0.5,r_{max}=2$
RandomFlip Randomly horizontally flip a image with probability of $p$. $p=0.5$
RandomCrop Randomly crop a image by a hight $h$ and a width $w$. $w=448,h=448$
GaussianBlur Randomly blur a image with probability of $p$. $p=0.5$
Transformation Description Parameter Setting
RandomRescale Rescale the image by $r$ times, $r$ randomly sampled from $r\sim U(r_{min},r_{max})$. $r_{min}=0.5,r_{max}=2$
RandomFlip Randomly horizontally flip a image with probability of $p$. $p=0.5$
RandomCrop Randomly crop a image by a hight $h$ and a width $w$. $w=448,h=448$
GaussianBlur Randomly blur a image with probability of $p$. $p=0.5$
OneOf Select one of the transformation in a transformation set $T$. $T=$ TransAppearance
Transformation Description Parameter Setting
Identity Returns the original image.
Autocontrast Maximizes the image contrast by setting the darkest (lightest) pixel to black (white).
Equalize Equalizes the image histogram.
RandSolarize Invert all pixels above a threshold value $T$. $T\in U(0,1)$
RandColor Adjust the color balance. $C=0$ returns a black&white image, $C=1$ returns the original image. $C\in U(0.05,0.95)$
RandContrast Adjust the contrast. $C=0$ returns a solid grey image, $C=1$ returns the original image. $C\in U(0.05,0.95)$
RandBrightness Adjust the brightness. $C=0$ returns a black image, $C=1$ returns the original image. $C\in U(0.05,0.95)$
RandSharpness Adjust the sharpness. $C=0$ returns a blurred image, $C=1$ returns the original image. $C\in U(0.05,0.95)$
RandPolarize Reduce each pixel to $C$ bits. $C\in U(4,8)$